OpenStudy (anonymous):

4 I 3a+4 I = 16a + 8

OpenStudy (anonymous):

413a+41=16a+8 subtract 16a from both sides: 397a+41=8 Then subtract 41 from both sides: 397a=-33 Then divide both sides by 397: a=-33/397

OpenStudy (anonymous):

I should have clarified that's the " l " between 4 and and 3 and again after 4 are absolute value sign's . 4 I 3a+4 I = 16a+8